# Learning to accelerate distributed ADMM using graph neural networks

### v1

@@ Abstract
Distributed optimization is fundamental in large-scale machine learning and control applications. Among existing methods, the Alternating Direction Method of Multipliers (ADMM) has gained popularity due to its strong convergence guarantees and suitability for decentralized computation. However, ADMM often suffers from slow convergence and sensitivity to hyperparameter choices. In this work, we show that distributed ADMM iterations can be naturally represented within the message-passing framework of graph neural networks (GNNs). Building on this connection, we propose to learn adaptive step sizes and communication weights by a graph neural network that predicts the hyperparameters based on the iterates. By unrolling ADMM for a fixed number of iterations, we train the network parameters end-to-end to minimize the final iterates error for a given problem class, while preserving the algorithm’s convergence properties. Numerical experiments demonstrate that our learned variant consistently improves convergence speed and solution quality compared to standard ADMM. The code is available at https://github.com/paulhausner/learning-distributed-admm.
@@ 1 Introduction
The emergence of large interconnected networked systems, where each agent in the network typically has access only to local private information, together with the recent explosion in data availability, has necessitated the development of efficient algorithms for distributed optimization across various domains such as machine learning, control, and signal processing. The alternating direction method of multipliers (ADMM) [22, 19] is well-suited for this task and has gained increased popularity due to its flexibility and scalability [8]. A key feature of ADMM is that it can be implemented in a distributed fashion, where the agents compute the iterations locally and in parallel. After computing their new iterates, the agents communicate the updates to their neighbors, and each node then updates its local variables by combining its own computations with the information received. While ADMM has a well-established convergence theory (see e.g. [16]), it is known to be highly dependent on the hyperparameters and may converge slowly to the global minimizer [42].
A promising avenue for addressing these limitations lies in recent advances in machine learning. In particular, the learning-to-optimize framework [12] aims to automate the design of faster algorithms by using data-driven techniques that enhance the convergence of iterative methods for a given problem class. Building on this idea, we adopt an unrolling scheme [30], where network parameters of our model are trained to minimize a loss function that reflects ADMM performance after a fixed number of iterations. In the context of distributed optimization problems, which consist of an interconnected network of agents, graph neural networks (GNNs) [7, 36] are a natural choice to parameterize the learning method as they can learn and reason about graph-structured data that can represent various problem domains such as interconnected network systems.
@@
Figure 1: Computation structure of iteration $k$ of the decentralized distributed ADMM algorithm computed as two message-passing steps in our proposed GNN (see Algorithm 2.2). The initial node features ${V}^{k}$ contain the previous ADMM iterates $(x^{k}_{i},y^{k}_{i},\lambda^{k}_{i})$ for each node $i$ and the non-zero edge features E{E} describe the weighted pairwise connection of nodes in the network which is fixed across iterations. After two message-passing steps – each consisting of a message (msg), an aggregation (agg), and an update (upd) part indicated by a colored box in the figure – the new iterates $(x^{k+1}_{i},y^{k+1}_{i},\lambda^{k+1}_{i})$ of all nodes in the network are the output of the GNN.
In particular, in this paper, we highlight that classical decentralized and distributed ADMM iterations [29] can be implemented as a graph network using the recently popularized message-passing framework [7] as we showcase in Figure 1 and further expand upon in Section 4.1. Throughout this paper, we explicitly exploit this connection and utilize GNNs to accelerate the distributed ADMM algorithm. The contributions of our paper can be summarized as follows:
  • We show the one-to-one correspondence of distributed ADMM iterations [29] and message-passing networks [21], a popular framework to build graph neural networks. This shows a strong alignment [40] between ADMM and GNNs which we exploit in this paper.
  • Based on these insights, we develop and evaluate different approaches for learning to accelerate distributed ADMM through learning the hyperparameters of the algorithm corresponding to graph-, node-, and edge-level tasks to learn the step size and communication matrix.
  • Finally, we propose a network architecture and loss functions to train the graph neural network in an end-to-end fashion by unrolling the ADMM iterations.
In numerical experiments, we demonstrate that our learning-based approaches reduces the distance of the iterates to the minimizer compared to the baseline method and can improve the convergence of ADMM even beyond the unrolling steps used during training. A concise summary of the notation used in this paper can be found in Appendix A.
@@ 2 Related work
The goal of learning-to-optimize (L2O) approaches is to accelerate the convergence of parameterized optimization problems using machine learning techniques [1, 12]. In particular, we focus on accelerating the ADMM algorithm for distributed optimization problems; for a detailed overview of the ADMM algorithm we refer to the survey by Boyd et al. [8].
@@
While L2O approaches can yield impressive results in practice [3], they typically lack convergence guarantees and may fail to generalize, especially on unseen problem instances or when the underlying data distribution shifts [1]. A popular approach to ensure convergence, which we also adopt, is to learn the hyperparameters of a classical optimization algorithm. Along those lines, Sambharya et al. [34] learn to warm start fixed-point algorithms, including ADMM, by unrolling the iterations and using automatic differentiation to compute the gradient through the iterations. In a follow-up work [33], they learn hyperparameters of fixed-point iterations. Similarly, Ichnowski et al. [28] use reinforcement learning to tune the hyperparameters of OSQP, an ADMM-based solver for convex quadratic programs. These approaches have also been combined to solve quadratic programming problems with distributed ADMM [35]. In contrast to this, we do not allow the presence of a central communication node. Further, we make no assumptions beyond the convexity of the objective function in our framework. A similar unrolling technique for ADMM iterations was successfully applied by Sjölund and Bånkestad [37] for nonnegative matrix factorization. Another approach by Noah and Shlezinger [31] learns the algorithms themselves directly through ADMM unrolling. We instead learn a network that predicts the step size hyperparameters of the optimization algorithm for a certain problem instance, either on a global or local level, increasing the flexibility of our learning method.
@@
Further, we propose a novel way to also learn the communication matrix. The predicted weighted Laplacian matrix can be seen as a form of learned preconditioning that stabilizes the iterates by improving problem conditioning [38]. Machine learning methods previously have been successfully applied to learn preconditioners for linear systems [24, 25], imaging problems [18], and gradient descent [20].
Since the communication network and the ADMM iterations can be represented as a graph, graph neural networks (GNNs) [7] are a natural fit for parameterizing the unrolled iterations. GNNs have been put to other L2O uses previously, as they can represent various optimization problems, such as quadratic programs [13], and can simulate the iterations of existing solvers such as interior point methods [32]. However, while Qian et al. [32] show that GNNs can, in principle, approximate the solution mapping by simulating the solver iterations, no convergence guarantees are obtained.
@@
Further, GNNs have also recently been applied to general distributed optimization. However, instead of starting with a learned algorithm and adding constraints inspired by theory [26], we formulate distributed ADMM as a GNN architecture operating on the communication network. Then, we parameterize the update functions with neural networks, allowing the GNN to learn on the given problem class, thereby replacing extensive manual hyperparameter tuning with a data-driven procedure.
@@ 3 Background
We start by briefly describing the distributed optimization problem setting and subsequently explain how ADMM iterations can be used to solve this problem. Finally, we give a short overview of the message-passing framework, which we later use to implement the ADMM iterations.
@@ 3.1 Distributed optimization
A distributed optimization problem consists of a network of $m$ interconnected agents, represented by the nodes $\mathcal{V}=\{1,2,\dots,m\}$ of a connected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. An edge $(i,j)\in\mathcal{E}$ symbolizes the opportunity for agent $i$ and $j$ to communicate. The agents jointly aim to solve the optimization problem
$\min_{x\in\mathbb{R}^{n}}\quad\sum_{i=1}^{m}f_{i}(x)\,$, (1)
where the local objective function $f_{i}\colon\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$ is only known to agent $i$ and its data parameterizing the objective cannot be communicated. Limited communication between agents can arise, for instance, due to limited networking capacity, as in edge networks [15], or due to privacy concerns [14]. In such scenarios, we must solve the problem in a distributed manner, decoupling the objective minimization and adding a consensus constraint:
$\min_{x_{1},x_{2},\dots x_{m}}\quad$ $\sum_{i=1}^{m}f_{i}(x_{i}),$ (2a)
s.t. $x_{1}=x_{2}=\dots=x_{m}\,$, (2b)
where each agent has its own local optimization variable $x_{i}\in\mathbb{R}^{n}$. Consensus among the agents can be reached by communicating these local solutions with their neighbors as defined by $\mathcal{G}$.
We assume that all $f_{i}$ are convex and that the objective (2a) is a proper, closed function. While there exist many different algorithms that solve this problem, we focus on distributed ADMM for peer-to-peer networks, i.e., the agents communicate directly with each other. In other words, we consider the decentralized setting, which is more general – and challenging – than having a central hub or intermediary.
@@ 3.2 ADMM
where $x\in\mathbb{R}^{p}$, $z\in\mathbb{R}^{q}$, $A\in\mathbb{R}^{r\times p}$, $B\in\mathbb{R}^{r\times q}$, and $c\in\mathbb{R}^{r}$. Originally proposed by Glowinski and Marroco [22] and Gabay and Mercier [19], ADMM combines features of dual ascent, allowing parallel updates, and the method of multipliers, which converges under mild assumptions [8].
@@
Following the approach proposed by Makhdoumi and Ozdaglar [29] we rewrite the distributed problem (2) to match the general ADMM formulation (3). In the rewritten form, the variable $x=[x_{i}]_{i=1}^{m}$ corresponds to the stacked individual variables of the agents and the function $f(x)=\sum_{i=1}^{m}f_{i}(x_{i})$ is the global objective. Further, the consensus constraint (2b) is enforced by the equivalent formulation $(P\otimes I_{n})x=0$, where $P$ is a $m\times m$ communication matrix satisfying $\text{null}(P)=\text{span}\{1_{m}\}$. This is in particular fulfilled by the (weighted) Laplacian $\mathcal{L}$ of a graph.
To allow parallel computation within ADMM, one can introduce an auxiliary variable $z$ and define $g$ as a convenient indicator function. Moreover, by exploiting the inherent symmetry of the optimization variables, the edge-based auxiliary variables $z_{ij}$ can be replaced with corresponding node-based variables $y_{i}$, while the original dual variables are reparameterized as related variables $\lambda_{i}$. This reparametrization reduces redundancy and leads to a more compact and efficient formulation [29].
@@
In the resulting distributed ADMM algorithm one iteration consists of updating all variables $x_{i}$, $y_{i}$, and $\lambda_{i}$ for each of the agents $i$. The updates can be executed for each agent $i$ in parallel and require only information about the local objective and previous iterates of its neighbors in the graph denoted by $\mathcal{N}(i)=\{j:(i,j)\in\mathcal{E}\}\subset\mathcal{V}$.  [^1] We do not allow for self-loops, so $i\notin\mathcal{N}(i)$. Further, since $\mathcal{G}$ is undirected $(i,j)=(j,i)\in\mathcal{E}$. The local solution candidates $x_{i}$ contain the solution to a regularized version of the local subproblem of agent $i$ and are computed as the minimizer to
$x_{i}^{k+1}=\arg\min_{x_{i}}\left(f_{i}(x_{i})\,+\hskip-7.11317pt\underset{j\in\mathcal{N}(i)\cup\{i\}}{\sum}\hskip-7.11317pt\left((\lambda_{j}^{k})^{T}(P_{ji}x_{i})+\frac{\alpha}{2}\|P_{ji}(x_{i}-x_{i}^{k})+y_{j}^{k}\|_{2}^{2}\right)\right)\,$, (4)
@@
Under the assumptions on the objective function from section 3.1, the ergodic average of the computed local solutions $x_{i}$ converges for every global step size $\alpha>0$ and any communication matrix $P$ with sublinear rate $\mathcal{O}(1/k)$ against the global solution. However, both the global step size $\alpha$ and the communication matrix $P$, affect the convergence speed in practice, and optimal choices appear to be instance-dependent [29]. The convergence results can be extended to the case where there exists not one but a sequence of step sizes [16], or even node-level step sizes $\alpha_{i}>0$ [6]. Moreover, the solution to (4) does not need to be exact to ensure convergence [16].
@@ 3.3 Graph neural networks
Graph neural networks (GNNs) learn a latent feature representation for each node in a graph based on some input features and the underlying graph topology. In this paper, we focus on the framework of message-passing neural networks (MPNNs) which update the node features of a graph by aggregating information from the neighboring nodes and edges [21]. Given edge weights $e_{ij}$ for $(i,j)\in\mathcal{E}$ represented by the weighted adjacency E{E} of the graph and initial node features $v_{i}^{0}$ for each node $i\in\mathcal{V}$, the latent node features $v_{i}^{l+1}$ are iteratively updated in each message-passing step $l$ by computing
@@
Here, the update and message functions can contain parameters that are learned during training, which are denoted by $\theta$. We denote the node feature matrix in layer $l$ with ${V}^{l}$. The aggregate function is typically permutation invariant, as there is no ordering of the graph neighbors, and can be implemented, for example, using the sum or mean function. This makes the overall architecture permutation equivariant. Thus, the ordering of the nodes does not affect the output of the network.
This framework can be easily extended to more general update and message functions, including global features and allowing nodes to send messages to themselves through self-loops [7].
@@ 4 Method
The key observation of our paper is that the distributed ADMM iterations introduced in Section 3.2 can be rewritten in the form of two message-passing steps of form (6). In the following, we explicitly establish the connection between the ADMM iterations and GNNs in Section 4.1. Further, we describe how the hyperparameters of the iterations can be parameterized using neural networks. In Section 4.2, we then describe how to train the parameters of this network using unrolled ADMM iterations.

### v2

@@ Abstract
[…]
@@ 1 Introduction
The emergence of large-scale interconnected networked systems has driven the need for efficient distributed optimization algorithms across machine learning, control, and signal processing. In these systems, each agent typically has access only to local information, yet must coordinate with neighbors to solve a global optimization problem. The alternating direction method of multipliers (ADMM) is well-suited for this task and has gained increased popularity due to its flexibility and scalability (boyd2011). A key feature of ADMM is that it can be implemented in a distributed fashion: agents perform computations locally and in parallel, then communicate updates to their neighbors and incorporate the received information into their local variables. While ADMM has a well-established convergence theory (eckstein1990admm), its convergence speed is highly sensitive to hyperparameter choices and can be slow in practice (zhang2019accelerating).
A promising avenue for addressing these limitations lies in recent advances in machine learning. In particular, the learning-to-optimize framework (chen2022learning) aims to automate the design of faster algorithms by using data-driven techniques that enhance the convergence of iterative methods for a given problem class. Building on this idea, we adopt an unrolling scheme (monga2021algorithm), where network parameters of our model are trained to minimize a loss function that reflects ADMM performance after a fixed number of iterations. In distributed optimization problems, which consist of an interconnected network of agents, graph neural networks (GNNs) (battaglia2018relational; scarselli2008graph) are a natural choice to parameterize the learning method as they can operate on graph-structured data that directly represents the underlying interconnected network.
@@
Figure 1: Computation structure of iteration $k$ of the decentralized distributed ADMM algorithm as two message-passing steps. The initial node features ${V}^{k}$ contain the previous ADMM iterates $(x^{k}_{i},y^{k}_{i},\lambda^{k}_{i})$ for each node $i$ and the non-zero edge features E{E} describe the weighted pairwise connections in the network which is fixed across iterations. After two message-passing steps – each consisting of a message (msg), an aggregation (agg), and an update (upd) part – the new iterates $(x^{k+1}_{i},y^{k+1}_{i},\lambda^{k+1}_{i})$ of all nodes in the network are the output of the GNN.
In this paper, we show that classical decentralized and distributed ADMM iterations (makh2017admm) can be implemented as a graph network using the message-passing framework (battaglia2018relational) as we showcase in Figure 1 and further expand upon in Section 4.1. We explicitly exploit this connection and utilize GNNs to accelerate distributed ADMM.
In summary, our contributions in this paper are threefold: First, we show the one-to-one correspondence between distributed ADMM iterations (makh2017admm) and message-passing networks (gilmer2017neural), establishing an explicit algorithmic alignment (velivckovic2021neural) between ADMM and GNNs. Second, based on these insights, we develop and evaluate different approaches for learning instance-specific hyperparameters, namely the communication matrix and local or global step sizes, to accelerate the distributed ADMM algorithm. These approaches correspond to graph-, node-, and edge-level learning tasks. Finally, we propose a network architecture and a loss function to train the graph neural network in an end-to-end fashion by unrolling the ADMM iterations, ultimately resulting in an automated instance-specific hyperparameter tuning of ADMM. In numerical experiments, we demonstrate that our learning-based approaches reduce the distance from the iterates to the minimizer compared to the baseline methods and can improve the convergence of ADMM even beyond the unrolling steps used during training.
@@ 2 Related work
The goal of learning-to-optimize (L2O) approaches is to accelerate the convergence of parameterized optimization problems using machine learning techniques (amos2023tutorial; chen2022learning). We focus on accelerating the ADMM algorithm for distributed optimization problems; for a detailed overview of the ADMM algorithm we refer to the survey by boyd2011.
@@
While L2O approaches can yield impressive results in practice (andrychowicz2016learning), they typically lack convergence guarantees and may fail to generalize, especially on unseen problem instances or when the underlying data distribution shifts (amos2023tutorial). A popular approach to ensure convergence, which we adopt, is to learn the hyperparameters of an existing optimization algorithm. Along those lines, sambharya2024learning learn to warm start fixed-point algorithms, including ADMM, by unrolling the iterations and using automatic differentiation to compute the gradient through the iterations. In a follow-up work, sambharya2024hyper learn hyperparameters of fixed-point iterations. Similarly, NEURIPS2021_afdec700 use reinforcement learning to tune the hyperparameters of OSQP, an ADMM-based solver for convex quadratic programs. These approaches have also been combined to solve quadratic programming problems with distributed ADMM (saravanos2024deep). In contrast, we do not assume the presence of a central communication node. Further, we make no assumptions beyond the convexity of the objective function. A similar unrolling technique for ADMM iterations was successfully applied by sjolund2022factorization for nonnegative matrix factorization. Another approach by noah2024distributed learns the algorithm parameters directly through ADMM unrolling on a fixed set of agents. We instead learn a network that predicts instance-specific hyperparameters of the optimization algorithm. biagioni2020learning use a recurrent neural network to directly predict the solution to ADMM which is used to warm-start the method.
@@
Additionally, we propose a novel way to learn the communication matrix. The predicted weighted Laplacian matrix can be seen as a form of learned preconditioning that stabilizes the iterates by improving problem conditioning (teixeira2015admm). Machine learning methods have previously been applied successfully to learn preconditioners for linear systems (hausner2024neural; hausner2025learning), imaging problems (fahy2024greedy), and gradient descent (gao2025gradient).
@@
GNNs have been used for other L2O purposes previously, as they can represent various optimization problems, such as quadratic programs (chen2024expressive; schmidtobreick2025warm), and can simulate the iterations of existing solvers such as interior point methods. However, while GNNs can, in principle, approximate the solution mapping by simulating the solver iterations, they provide no convergence guarantees (qian2024exploring). GNNs have also recently been applied to general distributed optimization. In contrast to approaches that start with a learned algorithm and add constraints inspired by theory (he2024mathematics), we formulate distributed ADMM as a GNN operating on the communication network. Then, we parameterize the iterations with neural networks, allowing the GNN to learn on the problem class, thereby replacing extensive manual hyperparameter tuning with a data-driven procedure.
@@ 3.1 Distributed optimization
A distributed optimization problem consists of a network of $m$ interconnected agents, represented by the nodes $\mathcal{V}=\{1,2,\dots,m\}$ of a connected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. An edge $(i,j)\in\mathcal{E}$ indicates that agent $i$ and $j$ can communicate. The agents jointly aim to minimize the sum of local objective functions $f_{i}$ where each $f_{i}\colon\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$ is only known to agent $i$ and the data parameterizing the objective cannot be communicated. Limited communication between agents can arise, for instance, due to limited networking capacity, as in edge networks (dal2023q), or due to privacy concerns (chen2024privacy). In such scenarios, we must solve the problem in a distributed manner, which requires decoupling the objective minimization and adding a consensus constraint:
$\min_{x_{1},x_{2},\dots x_{m}}\quad\sum_{i=1}^{m}f_{i}(x_{i})\quad\text{s.t.}\quad x_{1}=x_{2}=\dots=x_{m}\,$, (1)
where each agent has its own local optimization variable $x_{i}\in\mathbb{R}^{n}$. The agents can reach consensus by communicating their local solutions to their neighbors as defined by the graph $\mathcal{G}$.
We assume that all $f_{i}$ are convex and that the objective (1) is a proper, closed function. While there exist many different algorithms that solve this problem, we focus on distributed ADMM for peer-to-peer networks, i.e., the agents communicate directly with each other. In other words, we consider the decentralized setting, which is more general than having a central hub or intermediary.
@@ 3.2 ADMM
where $x\in\mathbb{R}^{p}$, $z\in\mathbb{R}^{q}$, $A\in\mathbb{R}^{r\times p}$, $B\in\mathbb{R}^{r\times q}$, and $c\in\mathbb{R}^{r}$. Originally proposed by Glowinski and Gabay, ADMM combines dual ascent, allowing parallel updates, and the method of multipliers, which converges under mild assumptions (boyd2011).
@@
Following the approach proposed by makh2017admm we rewrite the distributed problem (1) to match the general ADMM formulation (2). In the rewritten form, the variable $x=[x_{i}]_{i=1}^{m}$ corresponds to the stacked individual variables of the agents and the function $f(x)=\sum_{i=1}^{m}f_{i}(x_{i})$ is the global objective. Further, the consensus constraint in (1) is enforced by the equivalent formulation $(P\otimes I_{n})x=0$, where $P$ is a $m\times m$ communication matrix satisfying $\text{null}(P)=\text{span}\{1_{m}\}$. This is satisfied by the (weighted) Laplacian $\mathcal{L}$ of a graph.
To allow parallel computation within ADMM, one can introduce an auxiliary variable $z$ and define $g$ as a convenient indicator function. Moreover, by exploiting the inherent symmetry of the optimization variables, the edge-based auxiliary variables $z_{ij}$ can be replaced with corresponding node-based variables $y_{i}$, while the original dual variables are reparameterized as related variables $\lambda_{i}$. This reparametrization reduces redundancy and leads to a more compact and efficient formulation.
@@
Each iteration of the distributed ADMM algorithm updates all variables $x_{i}$, $y_{i}$, and $\lambda_{i}$ for each agent $i$. The updates can be executed in parallel and require only information about the local objective and previous iterates of the agent’s neighbors in the graph denoted by $\mathcal{N}(i)=\{j:(i,j)\in\mathcal{E}\}\subset\mathcal{V}$. [^1] We do not allow for self-loops, so $i\notin\mathcal{N}(i)$. Further, since $\mathcal{G}$ is undirected $(i,j)=(j,i)\in\mathcal{E}$. The local solution candidates $x_{i}$ is the solution to a regularized version of the local subproblem of agent $i$ and is computed as the minimizer of
$x_{i}^{k+1}=\arg\min_{x_{i}}f_{i}(x_{i})\,+\hskip-7.11317pt\underset{j\in\mathcal{N}(i)\cup\{i\}}{\sum}\hskip-7.11317pt\left((\lambda_{j}^{k})^{T}(P_{ji}x_{i})+\frac{\alpha}{2}\|P_{ji}(x_{i}-x_{i}^{k})+y_{j}^{k}\|_{2}^{2}\right)\,$, (3)
@@
Under the assumptions on the objective function from Section 3.1, the ergodic average of the computed local solutions $x_{i}$ converges for every global step size $\alpha>0$ and any communication matrix $P$ with sublinear rate $\mathcal{O}(1/k)$ to the global solution. However, both the global step size $\alpha$ and the communication matrix $P$, affect the convergence speed in practice, and optimal choices appear to be instance-dependent (makh2017admm). The convergence results can be extended to the case where there exists not one but a sequence of step sizes (eckstein1990admm), or even node-level step sizes $\alpha_{i}>0$ (barber2024node-level_alpha). Moreover, the solution to (3) does not need to be exact to ensure convergence (eckstein1990admm).
@@ 3.3 Graph neural networks
Graph neural networks (GNNs) learn a feature representation for each node in a graph based on input features and the underlying graph topology. In this paper, we focus on the framework of message-passing neural networks (MPNNs) which update the node features of a graph by aggregating information from the neighboring nodes and edges (gilmer2017neural). Given edge weights $e_{ij}$ for $(i,j)\in\mathcal{E}$ represented by the weighted adjacency E{E} and initial node features $v_{i}^{0}$ for each node $i\in\mathcal{V}$, the latent node features $v_{i}^{l+1}$ are iteratively updated in each message-passing step $l$ by computing
@@
Here, the update and message functions can contain parameters that are learned during training, which are denoted by $\theta$. We denote the node feature matrix in layer $l$ with ${V}^{l}$. The aggregate function is typically permutation invariant, as there is no ordering of the graph neighbors, and can be implemented, for example, using the sum or mean function. This makes the overall architecture permutation equivariant, so the ordering of the nodes does not affect the output of the network.
@@ 4 Method
Our key observation is that the distributed ADMM iterations introduced in Section 3.2 can be rewritten in the form of two message-passing steps of form (5). In Section 4.1, we explicitly establish the connection between the ADMM iterations and GNNs. Further, we describe how the hyperparameters of the iterations can be parameterized using neural networks. In Section 4.2, we then describe how to train the parameters of this network using unrolled ADMM iterations.